In a meat packaging plant, Machine A accounts for 60% of the plant's output while Machine B accounts for 40% of the plant's output. In total, 5% of the packages are improperly sealed. Also, 4% of the packages are from Machine A and are improperly sealed.

If a package selected at random is improperly sealed, what is the probability that it came from machine A?
Answer #1

By Baye's theorem,

P(A|I) = P(A)P(I|A) / P(I) = (0.60*0.04)/0.05 = 0.48